# Wildfire Risk Prediction: A Review

### v3

@@ 4.1.1 Fuel moisture content
Fuel Moisture Content (FMC) is defined as the ratio of the weight of water to the dry weight of the sample (Eq. 9)(Mbow et al., 2004; Chuvieco et al., 1999). FMC can be categorized into two types: Live Fuel Moisture Content (LFMC) and Dead Fuel Moisture Content (DFMC). It regulates the combustibility of the fuel (Yebra et al., 2013b) and is a significant factor influencing the spread rate of wildfires and the emissions produced during a fire (Fernandes et al., 2008; May et al., 2019).
$FMC=\frac{Weightofwaterinsample}{Dryweightofthesample}\times 100\%$ (9)
@@
In early studies, the live fuel moisture content (LFMC) was measured similarly to the dead fuel moisture content (DFMC) through field sampling and weighing methods, with interpolation or inversion used to estimate the entire study area (Bianchi and Defossé, 2015). While this method is highly accurate locally, it is time-consuming, labor-intensive, and not scalable to landscape, regional, or global levels (Yebra et al., 2013a). Like DFMC, LFMC can be estimated on a large scale in near real-time using meteorological parameters. This is because plant water regulation is influenced by environmental factors such as precipitation, soil moisture, and evapotranspiration (Bowyer and Danson, 2004). For example, the NFDRS used by the US Forest Service calculates LFMC using empirical formulas based on precipitation, temperature, and relative humidity (Bradshaw et al., 1984). Moreover, Castro et al. (2003) modeled the live fine fuel moisture content (LFFMC) of Cistus spp. in the Catalonia region of Spain using factors such as temperature, soil water availability, and atmospheric water content. Similarly, Ruffault et al. (2018) compared six drought indices’ ability to estimate LFMC across six different Mediterranean shrubs in southern France, finding that drought indices were effective only at stand scale, with indices simulating long-term drought dynamics (DC, RWC, and KBDI) providing better LFMC predictions. Viegas et al. (2001) used DC, BUI, and DMC from the FWI to estimate LFFMC in various plant species in central Portugal and Catalonia. Furthermore, Vinodkumar et al. (2021) successfully predicted LFMC across Australia using the lag relationship between soil moisture changes and vegetation moisture responses combined with regression analysis and physical models.
Meteorological data can also be combined with other datasets to estimate LFMC. For instance, Miller et al. (2023) proposed a method using a multimodal Multi-tempCNN model that incorporates multitemporal meteorological data, MODIS MCD43A4 BRDF-Adjusted Reflectance data, and climate zone, topography, and location information. This method enables large-scale LFMC estimation at medium resolution (500 m) over periods longer than two weeks. Similarly, McNorton and Di Giuseppe (2024) integrated soil moisture data and leaf area index (LAI) to estimate LFMC, validating the estimates using MODIS data. In the western United States, Wang and Quan (2023) used meteorological indicators, soil moisture data, and vegetation indices to estimate LFMC, finding that inversion accuracy varied with vegetation cover type.
The advantage of meteorological parameter-based LFMC estimation methods is the ability to generate near real-time LFMC estimates or predictions using data from weather stations, meteorological satellites, or numerical weather forecasting (Capps et al., 2021). However, LFMC is more challenging to estimate using meteorological parameters compared to DFMC. This is because plants actively regulate their water content through absorption and transpiration, with significant differences in physiological characteristics between plant species and a greater influence from phenological conditions and medium-term meteorological conditions (Pellizzaro et al., 2007; Ruffault et al., 2018; Jolly and Johnson, 2018).
Therefore, compared to LFMC estimation methods based on meteorological data or drought indices, most studies prefer using multi-source remote sensing data to estimate LFMC (Rao et al., 2020a; Capps et al., 2021; Quan et al., 2021a; Rodriguez-Jimenez et al., 2023). Remote sensing technology, as a low-cost means of large-scale, long-term dynamic Earth observation, also enables near real-time, high-resolution LFMC estimation in highly heterogeneous real-world scenarios (Quan et al., 2021b; Zhu et al., 2021; Verbesselt et al., 2002). The exponential growth of Earth observation missions (Ustin and Middleton, 2021) has provided new opportunities for LFMC estimation (Leite et al., 2024). Studies on LFMC estimation using multi-source satellite remote sensing primarily utilize multispectral reflectance data, thermal infrared data, and active and passive microwave data, combined with data-driven empirical or machine learning models, radiative transfer-based physical models, or hybrid models that integrate physical constraints with data-driven approaches (Yebra et al., 2013a; McNorton and Di Giuseppe, 2024).
The theoretical basis for LFMC inversion methods using multispectral reflectance data lies in the presence of water absorption bands in the near-infrared ( 750-1100 nm) and shortwave infrared ( 1100-2500 nm) spectral regions. Plant stress due to drought leads to changes in chlorophyll concentration and leaf internal structure, allowing the relationship between vegetation spectral reflectance (i.e., spectral characteristics) and LFMC to be established (Knipling, 1970; Kriedemann and Barrs, 1983; Bowyer and Danson, 2004; Yebra et al., 2013a; Sow et al., 2013; Zacharakis and Tsihrintzis, 2023b; Leite et al., 2024). Thus, LFMC can be inverted by simulating the relationship between spectral characteristics under different water content levels and observed spectral characteristics (Hao and Qu, 2007; Yebra et al., 2008, 2013a; Marino et al., 2020; Quan et al., 2021a). While this physical method is generally more robust, it is also more complex in terms of parameterization (Yebra et al., 2013a). Alternatively, empirical methods based on spectral reflectance or spectral indices, including the normalized difference vegetation index (NDVI), enhanced vegetation index (EVI), soil-adjusted vegetation index (SAVI), normalized difference infrared index (NDII), water index (WI), normalized difference water index (NDWI), visible atmospheric resistance index (VARI), Leaf Water Content Index (LWCI) (Hunt Jr et al., 1987), Global Vegetation Moisture Index (GVMI), and Shortwave Infrared Water Stress Index (SIWSI) (Fensholt and Sandholt, 2003), are simpler and only require fitting the relationship between in situ LFMC measurements and spectral reflectance characteristics. These methods can perform as well as or better than physical models at local scales and across specific vegetation types (Cunill Camprubi et al., 2022).
For example, Chuvieco et al. (2002) compared the ability of shortwave infrared bands, NDVI, NDII, LWCI, the wetness and greenness components of the Tasseled Cap transformation, and spectral derivatives to model LFMC in Mediterranean grasslands and shrubs. They found that shortwave infrared bands and indices using these bands performed better due to the high correlation between shortwave infrared and LFMC, attributed to water absorption in leaves. Near-infrared and red bands also showed good correlation with LFMC, particularly for herbaceous plants, where drought stress led to changes in chlorophyll content and structure. The study recommended further use of thermal bands and meteorological data to improve modeling accuracy. Additionally, Myoung et al. (2018) used Pearson correlation analysis to compare the correlations between 16-day NDVI and EVI products (MOD13Q1 and MYD13Q1) and NDWI, NDII, and VIARI calculated from MOD09A1 reflectance data and in situ LFMC measurements from the National Fuel Moisture Database. The study ultimately selected EVI and introduced daily minimum temperature data to establish an empirical model for estimating LFMC in southern California. This choice was because EVI effectively captured changes in chlorophyll concentration related to water content, showing better response than other vegetation indices and shortwave infrared bands under water stress, particularly for Mediterranean vegetation. Introducing temperature data improved the model’s ability to capture excessive water loss during extreme heat, correcting overestimates of minimum LFMC. Similarly, García et al. (2020) found that EVI performed better than other ratio vegetation indices, such as NDII, NDVI, and VARI, for LFMC modeling in Eastern California and Oregon.
Marino et al. (2020) used NDVI, NDII, GVMI, NDWI, EVI, SAVI, VARI, and $VI_{green}$ from MCD43A4 and MOD09GA products, as well as the same indices calculated from Sentinel-2 data, to estimate the LFMC of Mediterranean shrubs. The main difference between MOD09GA and MCD43A4 and the MOD13Q1 and MYD13Q1 products used by Myoung et al. (2018) is that the former are daily data, which do not smooth out short-term LFMC changes. This feature is more conducive to capturing abnormal LFMC values under extreme conditions and to estimating LFMC for shrubs and grasslands that are more sensitive to short-term environmental changes. Similar to Myoung et al. (2018), Marino et al. (2020) also found that using vegetation indices to estimate LFMC tends to overestimate lower LFMC values. However, Marino et al. (2020) found that VARI had the highest accuracy in LFMC modeling compared to other vegetation indices. Combining VARI with NDVI can further improve the accuracy of the model. This is because VARI and NDVI are based on visible (blue, green, red) and near-infrared reflectance data, which are suitable for measuring changes in greenness and LAI related to leaf drying and can indirectly reflect leaf water content. Similarly, Stow et al. (2006); Peterson et al. (2008); Caccamo et al. (2011) found that VARI performed better than other vegetation indices or reflectance bands in estimating LFMC for shrubs in Southern California, Mediterranean climate regions, and southeastern Australia. Although these studies did not show that any particular vegetation index is the best choice in all scenarios, they emphasized the importance of chlorophyll (or greenness) response to water stress.
Water availability is a key parameter for plant transpiration. Therefore, changes in LFMC affect not only reflectance spectral bands but also canopy temperature (ST). When plants dry out, transpiration and latent heat transfer decrease, which in turn increases sensible heat (Kozlowski et al., 1990; Yebra et al., 2013a). As a result, introducing thermal infrared bands (2500-14000 nm) into LFMC estimation can improve model performance under certain conditions. It is worth noting that in most studies, land surface temperature (LST) is used as a proxy for ST. Chuvieco et al. (2004) used NDVI and ST data from the AVHRR sensor to estimate LFMC for grasslands and shrubs in Cabañeros National Park, central Spain, based on a linear regression model. The study found that NDVI and ST were positively and negatively correlated with LFMC, respectively, and that the introduction of ST significantly improved the model’s performance. Cunill Camprubi et al. (2022) combined MODIS spectral bands, vegetation indices, land surface temperature, and day of the year as predictors, using a random forest algorithm to estimate LFMC in the Mediterranean basin wildlands. They found that the combination of LST and day of the year had the highest importance, and that when the temperature exceeded 20°C, LST was negatively correlated with LFMC. However, in LFMC estimation experiments based on MODIS LST products, LST did not show a significant advantage over other factors. For example, Sow et al. (2013) compared the performance of the NDVI-to-LST ratio and GVMI, WI, NDWI, NDII, and MSI in estimating LFMC and equivalent water thickness (EWT) in three study areas in Senegal, finding that the NDVI-to-LST ratio had no significant advantage in correlating with LFMC and EWT compared to other vegetation indices. It is worth noting that this study did not compare the performance of using NDVI alone to model LFMC or EWT. (McCandless et al., 2020) used machine learning methods to model LFMC across the continental United States using MODIS reflectance data, meteorological data, and terrain data. The random forest method was used to evaluate the importance of various predictors, finding that the importance of LST was only 6
Although spectral features can model canopy LFMC, they are more sensitive to dry matter such as canopy structure, resulting in lower robustness (Rao et al., 2020b). Additionally, these spectral features couple dry matter and water content information, so when both change due to water stress, there can be issues with estimating LFMC using vegetation indices and other spectral features (Danson and Bowyer, 2004; Yebra et al., 2013b). Moreover, spectral features are affected by lighting conditions and clouds or rain, making it difficult to provide full-time or near-real-time large-scale, long-term series LFMC mapping (Rao et al., 2020b). Compared to spectral reflectance, microwave remote sensing is more capable of penetrating clouds and is more sensitive to water content (Rao et al., 2020b). For example, active microwave sensors are sensitive to vegetation structure or dielectric properties (including vegetation moisture) through backscatter signals, and their penetration capability can be used to model soil moisture, thus monitoring vegetation water stress (Dorigo et al., 2017). Passive microwave sensors can also be used to estimate LFMC through soil moisture time lags or vegetation optical depth (VOD) (Lu and Wei, 2021a; Forkel et al., 2023). Commonly used microwave sensors primarily operate in the P-band ($\approx$ 70 cm), L-band ($\approx$ 23 cm), C-band and X-band ($\approx$ 3 cm), and Ku-band ($\approx$ 1.6-2.5 cm). Among them, the backscatter signal of P and L bands is primarily coupled with backscatter from the soil and lower canopy, while the backscatter signal of C, X, and Ku bands mainly comes from the interaction between microwave radiation and the upper canopy.
Tanase et al. (2015) used vertical-vertical (VV), vertical-horizontal (VH), horizontal-vertical (HV), and horizontal-horizontal (HH) polarization L-band radar backscatter signals to model the relationship between backscatter intensity, polarization decomposition signals, radar vegetation index (RVI) (Kim et al., 2012), and canopy height with LFMC and EWT. Experiments in Australian forests demonstrated that the modeling capability of L-band radar data was superior to spectral reflectance data. The penetration of L-band data enables modeling of the lower canopy and surface fine fuel moisture content. However, when the forest is high and dense, and both lower canopy and soil moisture are low, there will be a large error in estimating LFMC for the lower canopy. Wang et al. (2019) explored the performance of C-band Sentinel-1A (5.4 GHz) VV and VH polarization data in estimating LFMC in Texas forests. The model coupled an empirical bare soil backscatter model with the Water Cloud Model (WCM) (De Roo et al., 2001) to establish an LFMC lookup table, directly estimating LFMC, and compared this LFMC with results obtained from Landsat 8 OLI spectral reflectance data using the empirical partial least squares (PLSR) method and measured LFMC data. The experiment showed that the accuracy of LFMC inversion based on Sentinel-1A data was higher than that based on Landsat 8 OLI data and that it could reproduce the temporal variation trend of LFMC. Lu and Wei (2021b) used multi-source microwave soil moisture products ECV_SM and SMAP L-band Level 4 surface and root zone soil moisture data, using a robust linear regression model with a time lag to assess LFMC across the continental United States. The study showed that LFMC was most correlated with soil moisture conditions 60 days before LFMC sampling in most regions.
Considering the influence of soil moisture content and canopy dry matter on radar backscatter, reflectance, or emission signal intensity, some studies have integrated SAR and multispectral reflectance data, along with vegetation cover, canopy height, and soil moisture data, to estimate LFMC. For instance, Jia et al. (2019) employed soil moisture data obtained from the SMAP L-band radiometer, combined with VARI calculated from MODIS MCD43A4 reflectance data, growing degree days (GDD), and precipitation, to estimate LFMC in Southern California using a regression model. This study found that using cumulative soil moisture data combined with cumulative growing degree days yielded better performance in LFMC estimation compared to using MODIS VARI combined with cumulative growing degree days. The superior performance is attributed to a more accurate representation of the water cycle processes between vegetation and soil.
Rao et al. (2020b) designed a knowledge-assisted neural network to estimate LFMC in the western United States, utilizing Sentinel-1 VV and VH polarized SAR backscatter data, Landsat-8 red, green, blue, near-infrared (NIR), and shortwave infrared (SWIR) optical reflectance data, along with NDVI, NDWI, and near-infrared spectral indices derived from the reflectance data. To separate the contributions of vegetation and soil moisture to the backscatter signal, the study proposed a knowledge-based method and a data-driven approach. The former method leverages the differential sensitivity of VV and VH polarizations to soil and vegetation moisture, using the difference between VV and VH polarizations to separate the contributions of soil and vegetation to backscatter intensity. The latter approach involves training the deep learning model to learn the separation of vegetation moisture and soil moisture contributions by incorporating soil type as an input. The study also established a ratio of backscatter signals to spectral reflectance, utilizing the high water sensitivity of backscatter signals and the high spectral reflectance of dry matter, alongside the aforementioned vegetation indices to separate the contributions of dry matter and water content to LFMC estimation.
Similarly, Xie et al. (2022a) employed Sentinel-1 data along with MODIS MCD43A4 nadir BRDF-adjusted reflectance (NBAR) data, original band reflectance data from Landsat-8 OLI (red, green, blue, NIR, and SWIR channels), and vegetation indices such as NDVI, NIRV, and NDWI. The study used canopy height and land cover type as auxiliary data to estimate LFMC in the western United States. These data were interpolated to unify spatial and temporal resolutions before being input into an integrated model, where LFMC was predicted through end-to-end training.
Furthermore, Forkel et al. (2023) achieved global LFMC estimation by combining Ku-band, X-band, and C-band VOD data derived from passive microwave technology with MODIS LAI, AVHRR vegetation cover data, LULC, and meteorological and topographical data. This study demonstrated that the combination of Ku-VOD and LAI as predictors significantly improved LFMC estimation performance. Conversely, Chaparro et al. (2024) conducted LFMC inversion experiments in the western United States, showing that when a model effectively separates vegetation structure and dry matter information, the accuracy of LFMC estimation based on VOD is independent of canopy height, land cover type, and radar backscatter signals.
While the aforementioned LFMC estimates based on multispectral and microwave remote sensing data have achieved satisfactory accuracy, it is noteworthy that some studies suggest that the inclusion of microwave data does not necessarily enhance LFMC inversion accuracy. For example, Tanase et al. (2022) explored the estimation of LFMC in Mediterranean vegetation using C-band Sentinel-1 and Sentinel-2 multispectral data combined with canopy height, vegetation cover, and topographical data. The study found that incorporating radar data did not improve model performance and thus only recommended its use in areas with significant cloud cover.
Research has also been conducted on the effectiveness of different data sources for LFMC estimation. For instance, Fan et al. (2018) compared the performance of LFMC estimation in southern France using ECV_SM, root zone soil moisture from ECV_SM, microwave polarization difference index (MPDI) calculated from C, X, Ku, K, and Ka band passive microwave data at 6.9, 10.7, 18.7, 23.8, and 36.5 GHz, VOD derived from AMSR-E C and X band data, and vegetation indices such as NDVI, SAVI, VARI, NDWI, NDII, and GVMI calculated from MODIS data. The MPDI and VOD were averaged over 15 days prior to in-situ measurements. The study found that the optical vegetation indices VARI and SAVI outperformed the microwave-derived indices and other optical indices. Among the microwave indices, X-band VOD showed the highest consistency with in-situ measured LFMC.
@@ 4.1.2 Fuel load
The methods for estimating fuel load are similar to those for estimating FMC, including direct sampling and the use of physical models or multispectral and microwave remote sensing data. Direct measurement involves collecting field samples, weighing, drying them, and then estimating biomass in the area, which is considered the most accurate assessment method and can provide a data foundation for physical or empirical model evaluations. First, fuel load estimation methods based on ground-measured data combined with vegetation or forest growth models can predict or estimate fuel load by forecasting future vegetation growth, forest structure, and more. For example, Lee et al. (2022) used a Weibull function and mortality model based on forest inventory data to predict the fuel load of Korean red pine (Pinus densiflora) for the next 20 years in Korea. Meanwhile, Nolan et al. (2022) modeled above-ground biomass in the shrub layer of eucalyptus forests in southeastern Australia using allometric equations, considering it as fuel load. Furthermore, Li and He (2022) developed a semi-empirical fuel load estimation model for the Krycklan region in northwestern Sweden by combining L-band ALOS PALSAR and Landsat 7 ETM+ reflectance data, along with NDII, NDVI, NDTI, and RVI vegetation indices, using vegetation cover and land cover type as auxiliary data. The study used the Water Cloud Model (WCM) to simulate forest scattering and corrected the influence of vegetation cover on the simulated scattering signal using reflectance data.
@@
In addition, the more recent NASA Global Ecosystem Dynamics Investigation (GEDI) full-waveform LiDAR sensor has been widely used for large-scale above-ground biomass estimation and vegetation canopy structure analysis (Myroniuk et al., 2023; Liang et al., 2023; Aragoneses et al., 2024; Guerra-Hernández et al., 2024), aiding wildfire simulation, assessment of wildfire impacts on vegetation, and understanding wildfire severity. However, only a few studies have directly used GEDI LiDAR data to estimate fuel load. For example, Leite et al. (2022) used unmanned aerial vehicle (UAV) simulations and GEDI data to study the ability to stratify fuel load in grasslands, savannas, and forests within tropical savannas in Brazil. However, the study found that GEDI performed poorly in estimating fuel load for ground and low-lying vegetation, and was significantly affected by terrain.

### v4

@@ 4.1.1 Fuel moisture content
FMC is defined as the ratio of the weight of water (WoW) to the dry weight (DW) of the sample (Eq. 7)(Mbow et al., 2004; Chuvieco et al., 1999). FMC can be categorized into two types: Live Fuel Moisture Content (LFMC) and Dead Fuel Moisture Content (DFMC). It regulates the combustibility of the fuel (Yebra et al., 2013b) and is a significant factor influencing the spread rate of wildfires and the emissions produced during a fire (Fernandes et al., 2008; May et al., 2019).
$FMC=\frac{WoW}{DW}\times 100\%$ (7)
@@
In early research, LFMC was measured analogously to DFMC through field sampling and weighing, with broader area estimates obtained via interpolation or inversion techniques (Bianchi and Defossé, 2015). While locally precise, this method is neither scalable nor efficient for landscape or global applications due to its time-consuming and labor-intensive nature (Yebra et al., 2013a). Additionally, meteorological parameter-based LFMC estimation benefits from near real-time predictions but faces challenges due to the active regulation of water content by plants and their phenological differences (Pellizzaro et al., 2007; Ruffault et al., 2018).
As such, many studies favor multi-source geospatial data for large-scale and near real-time LFMC estimation, utilizing the growing capabilities of Earth observation technologies for high-resolution data (Rao et al., 2020a; Capps et al., 2021). These contemporary approaches utilize environmental variables such as precipitation and soil moisture, along with spectral, thermal infrared, and microwave data, often in combination with machine learning algorithms, to enhance prediction accuracy (Bowyer and Danson, 2004; Yebra et al., 2013a).
LFMC inversion methods based on multispectral reflectance data exploit water absorption bands in the near-infrared and shortwave infrared spectral regions. Drought stress leads to changes in chlorophyll concentration and leaf structure, enabling LFMC estimation through spectral reflectance relationships (Knipling, 1970; Yebra et al., 2013a). Physical models simulate these relationships, but they are complex; empirical methods using spectral indices like NDVI, EVI, and NDWI provide simpler alternatives and perform well at local scales (Hunt Jr et al., 1987; Fensholt and Sandholt, 2003). Studies have demonstrated the effectiveness of shortwave infrared bands and vegetation indices for LFMC estimation, though improvements through integrating meteorological data have also been suggested (Chuvieco et al., 2002; Myoung et al., 2018).
Thermal infrared bands, such as land surface temperature (LST), are also incorporated to enhance LFMC estimation since canopy temperature correlates with plant water stress (Chuvieco et al., 2004). Machine learning methods have further combined spectral and thermal data to model LFMC, revealing varying importance levels among predictors (McCandless et al., 2020).
Remote sensing methods often face limitations in lighting conditions and cloud coverage, prompting the use of microwave sensors for LFMC estimation due to their ability to penetrate clouds and sensitivity to water content (Rao et al., 2020b). Both active and passive microwave sensors have been employed, with studies showing promising results for vegetation monitoring and LFMC estimation (Dorigo et al., 2017; Lu and Wei, 2021). For example, integrating SAR and multispectral data with vegetation indices improves LFMC predictions, as shown in studies using Sentinel-1 and Landsat-8 data (Jia et al., 2019; Rao et al., 2020b).
Overall, advancements in combining multi-source data and modeling techniques present significant opportunities for accurate LFMC estimation. However, it is important to note that due to the complexity of real-world scenarios, LFMC estimation still faces significant challenges. These challenges include separating the effects of vegetation dry matter, soil moisture, and atmospheric and canopy structure interference on backscatter and optical reflectance and emissivity. Moreover, LFMC ground truth data remain scarce and sparsely distributed, hindering the development of data-driven algorithms.
@@ 4.1.2 Fuel load
The methods for estimating fuel load are similar to those for estimating FMC, including direct sampling and the use of physical models or multispectral and microwave remote sensing data. Direct measurement involves collecting field samples, weighing, drying them, and then estimating biomass in the area, which is considered the most accurate assessment method and can provide a data foundation for physical or empirical model evaluations. First, fuel load estimation methods based on ground-measured data combined with vegetation or forest growth models can predict or estimate fuel load by forecasting future vegetation growth, forest structure, and more. For example, Lee et al. (2022) used a Weibull function and mortality model based on forest inventory data to predict the fuel load of Korean red pine (Pinus densiflora) for the next 20 years in Korea. Meanwhile, Nolan et al. (2022) modeled above-ground biomass in the shrub layer of eucalyptus forests in southeastern Australia using allometric equations, considering it as fuel load.
@@
In addition, the more recent NASA Global Ecosystem Dynamics Investigation (GEDI) full-waveform LiDAR sensor has been widely used for large-scale above-ground biomass estimation and vegetation canopy structure analysis (Myroniuk et al., 2023; Liang et al., 2023; Aragoneses et al., 2024; Guerra-Hernández et al., 2024), aiding wildfire simulation, assessment of wildfire impacts on vegetation, and understanding wildfire severity. However, only a few studies have directly used GEDI LiDAR data to estimate fuel load. For example, Leite et al. (2022) used unmanned aerial vehicle (UAV) simulations and GEDI data to study the ability to stratify fuel load in grasslands, savannas, and forests within tropical savannas in Brazil. The study found that GEDI performed poorly in estimating fuel load for ground and low-lying vegetation, and was significantly affected by terrain.